Kenneth's father is 36 years older than he is. In 7 years' time, his father will be double Kenneth's age. How old is Kenneth's father now?
Answer by greenestamps(13200)



A formal algebraic solution....

Let x be the age of Kenneth's father
Then x-36 is Kenneth's age

x+7 = father's age in 7 years
(x-36)+7 = x-29 = Kenneth's age in 7 years

In 7 years, the father's age will be twice Kenneth's age:

x+7=2(x-29)
x+7=2x-58
65=x

ANSWER: 65

CHECK: x+7=65+7=72; x-29=65-29=36; 72=2*36

Informally (using logical reasoning and simple mental arithmetic, which is good exercise)....

The difference in their ages is 36; when the father is twice as old, their ages will be 72 and 36. That will be 7 years from now, so the father's age now is 72-7 = 65